OpenStudy (anonymous):

Calculate the H30+ concentration in 6M NaOH.

OpenStudy (anonymous):

1.0x10^-14/6=1.67*10^-15 M H3O+ Kw=1.0x10^-14=[H3O+][OH-] [H3O+]=Kw/[OH-]=1.0x10^-14/[OH-]

OpenStudy (anonymous):

10^-8